-4t - 2(7 - 3t)= -3t + 23 i can;t figure this out
Found 2 solutions by terbear1963, funmath:
Answer by terbear1963(43)

FIRST YOU MUST GET YOUR T ON ONE SIDE AND YOUR NUMBERS ON ONE SIDE!!!! LET US GET T ON THE LEFT AND THE NUMBERS ON THE RIGHT FOR THIS PROBLEM. -4T-2(7-3T)= -3T+23 ---> -4T-14+6T=-3T+23 ---> 2T-14=-3T+23 ---> 5T=37 ---> T=37/5 WHEN YOU PLUG 37/5 IN YOUR EQUATION FOR T YOU WILL FIND, FINALLY, THAT YOUR PROBLEM WORKS OUT AND IS EQUAL!!! I HOPE THIS HELPS YOU OUT!? YOUR MATH FRIEND, TERRY (cats7051@yahoo.com)

Answer by funmath(2933)

-4t - 2(7 - 3t)= -3t + 23
-4t-2(7)-2(-3t)=-3t+23
-4t-14+6t=-3t+23
2t-14=-3t+23
2t+3t-14=-3t+3t+23
5t-14=23
5t-14+14=23+14
5t=37
5t/5=37/5
t=37/5
Happy Calculating!!!